Suppose you just received a shipment of thirteen televisions. Three of the televisions are defective. If two televisions are randomly​ selected, compute the probability that both televisions work. What is the probability at least one of the two televisions does not​ work?
Answer #1

In 13 TVs, 3 are defective

a)

2 TVs from 13 can be selected in 13C2 ways

2 non defective TVs from 11 can be selected in 11C2 ways

required probability = 11C2/13C2 = 55/78 = 0.7051

b)

Probability that atleast one of 2 TVs does not work = 1 - P(both work)

= 1- 0.7051

= 0.2949